# FailSafe: Reasoning and Recovery from Failures in Vision-Language-Action Models

### v1

@@ Abstract
Recent advances in robotic manipulation have integrated low-level robotic control into Vision-Language Models (VLMs), extending them into Vision-Language-Action (VLA) models. Although state-of-the-art VLAs achieve strong performance in downstream robotic applications, supported by large-scale crowd-sourced robot training data, they still inevitably encounter failures during execution. Enabling robots to reason about and recover from unpredictable and abrupt failures remains a critical challenge. Existing robotic manipulation datasets, collected in either simulation or the real world, primarily provide only ground-truth trajectories, leaving robots unable to recover once failures occur. Moreover, the few datasets that address failure detection typically offer only textual explanations, which are difficult to utilize directly in VLA models. To address this gap, we introduce FailSafe, a novel failure generation and recovery system that automatically produces diverse failure cases paired with executable recovery actions. FailSafe can be seamlessly applied to any manipulation task in any simulator, enabling scalable creation of failure–action data. To demonstrate its effectiveness, we fine-tune LLaVa-OneVision-7B (LLaVa-OV-7B) to build FailSafe-VLM. Experimental results show that FailSafe-VLM successfully helps robotic arm detect and recover from potential failures, improving the performance of three state-of-the-art VLA models ($\pi_{o}$-FAST, OpenVLA, OpenVLA-OFT) by up to 22.6% on average across several tasks in Maniskill. Furthermore, FailSafe-VLM could generalize across different spatial configurations, camera viewpoints, and robotic embodiments. We plan to release the FailSafe code to the community. Project Page: https://jimntu.github.io/FailSafe/
@@ I INTRODUCTION
Figure 1: An illustration of the FailSafe pipeline generating failure scenarios and corresponding executable recovery actions (above). Leveraging these, FailSafe enables FailSafe-VLM (below) to detect and recover from robot failures, while generalizing across different spatial configurations, viewing angles, and embodiments.
@@
Furthermore, unlike prior work that frames failure reasoning merely as textual explanations [22, 19, 20], FailSafe collects executable recovery actions that can be directly applied by robots in real time. To ensure accuracy and robustness, each recovery action is validated through a rigorous sanity check, confirming that it effectively resolves the failure scenario and eventually leads to successful task completion. By incorporating multi-view images and task instructions, this pipeline produces a customized, accurate, and explainable FailSafe dataset for robotic manipulation that would be prohibitively difficult to obtain through real-world data collection.
Experimental results further highlight the effectiveness of FailSafe in VLA deployments. We fine-tune LLaVa-OV-7B [23] to obtain FailSafe-VLM, which demonstrates strong failure reasoning capabilities that generalize across diverse spatial configurations, camera viewpoints, and robotic embodiments. FailSafe-VLM significantly outperforms state-of-the-art VLMs, such as GPT-4o [24] and Gemini-2.5-flash [25], in detecting failures and generating accurate recovery actions on unseen failure trajectories. Furthermore, when serving as an external assistant to VLA models, FailSafe-VLM identifies and corrects potential failures in real time during robot execution, leading to an average performance improvement of up to 22.6% on ManiSkill tasks [26] compared to three VLA baselines without FailSafe-VLM. Finally, FailSafe-VLM also proves effective when evaluated on a previously unseen robot arm, demonstrating its ability to generalize across different embodiments.
@@
  • We show that FailSafe dataset enables existing VLMs to reason about failures and significantly improves the performance of VLA models across diverse viewpoints, spatial configurations, and robotic embodiments.
@@ II-A Vision-Language-Action model
Figure 2: Top: Overall pipeline of FailSafe, which includes the autonomous generation of failure trajectories (I) and collection of delta recovery action (II). Failure-Action data pairs are passed to the next step only after a sanity check (III) ensures the effectiveness of recovery action. Bottom: The FailSafe dataset (IV) is then used to fine-tune FailSafe-VLM, which is able to help robotic arms recover from failure cases (V).
@@ III FailSafe
We propose a novel pipeline, FailSafe, to systematically collect failure cases and corresponding recovery actions within ManiSkill [26]. While developed in ManiSkill, this pipeline is designed to generalize across a broader range of simulators that support motion planning. To ensure the reliability of the collected data, we incorporate a sanity check mechanism that validates each failure–action pair before inclusion in the dataset. This step is crucial to guarantee both the diversity and effectiveness of the recovery actions. Details of the failure generation, action collection and sanity check mechanism are provided in Section III-A, Section III-B and Section III-C, respectively.
FailSafe dataset is subsequently used to fine-tune LLaVa-OV-7B [23], yielding FailSafe-VLM, which is an expert vision–language model capable of assisting VLA models in reasoning about and recovering from failures during robotic manipulation. The dataset format and the fine-tuning process for obtaining FailSafe-VLM are described in Section III-D and Section III-E, respectively.
@@ III-B Action Collection
III-C Sanity Check
After getting the candidate corrective actions, they need to go through rigorous sanity check to be included into the final FailSafe dataset. Importantly, the sanity check is designed to ensure that the collected recovery actions $\Delta A$ are both robust in correcting failure cases and effective when applied at any timestep where failures could potentially occur in the immediate future.
In sanity check, FailSafe replays the trajectory by incorporating the two poses ($P_{d}$ and $P_{c}$) collected earlier. Specifically, the motion planner first moves the robotic arm to the deviated pose $P_{d}$, then to the corrective pose $P_{c}$, and finally continues with the subsequent poses of the task (i.e., $A\rightarrow P_{d}\rightarrow P_{c}\rightarrow B\rightarrow C\rightarrow D$). If the originally failed operation is successfully completed after applying the corrective pose $P_{c}$, the corrective action $\Delta A$ along with the information such as image observation and failure types collected during the failure generation phase is added to the dataset. This process ensures both the reliability and diversity of the collected recovery actions, thereby enhancing FailSafe-VLM’s ability to help VLA models recover from failure scenarios in manipulation tasks.
@@ IV Experiments
In this section, we design a series of experiments to evaluate the failure reasoning and recovery capabilities of FailSafe-VLM in Framka Emika Panda robot arm. First, we roll out test seeds in Maniskill to generate unseen failure and success trajectories with new spatial configurations, and compare FailSafe-VLM with three other VLMs to assess the accuracy of predicted failure types and recovery actions. Second, we integrate FailSafe-VLM with three state-of-the-art VLA models to evaluate whether it can help them recover from failures and improve overall performance. We also step further to investigate the cross-embodiment generalization capability of FailSafe-VLM. Finally, a qualitative analysis is conducted to visualize the end-effector pose when the VLA model is integrated with FailSafe-VLM.
@@ IV-B Performance on VLA
TABLE III: Success rates of VLA models with and without FailSafe-VLM on three ManiSkill tasks using the Franka Emika Panda robot arm; the final row reports results on xArm 6. The experimental camera view matches the training view of the VLA models but is novel to FailSafe-VLM, and improvements are reported under this setting.
The experimental results in Table III show that all three VLA models yield average performance improvements with the inclusion of FailSafe-VLM. In particular, OpenVLA improves by more than 20%, with an average gain of 22.6% across the three tasks. For OpenVLA-OFT and $\pi_{o}$-FAST, whose baselines are already relatively strong, failure reasoning still pushes the boundary, increasing success rates by 8.0% and 4.0%, respectively. We further evaluate whether FailSafe-VLM can assist VLA models in an embodiment unseen during training. Similarly, we also collect 1,000 trajectories per task on xArm 6 to fine-tune OpenVLA-OFT, while reusing the same FailSafe-VLM checkpoint trained exclusively on Franka Emika Panda robots. Although FailSafe-VLM is never trained on xArm 6 data, the failure scenarios and trajectories generated by FailSafe are independent of the specific robotic embodiment. Consequently, FailSafe-VLM exhibits cross-embodiment generalization by boosting the performance of the stack cube task on xArm 6 from 56% to 76% without degrading performance on other tasks.
For seeds recovered by FailSafe-VLM, we observe helpful nudges when the robotic arm becomes stuck or is about to fail. We would like to highlight that FailSafe-VLM should be viewed as an assistant for failure reasoning and recovery, rather than as a replacement for VLA models. Overall, these results demonstrate that the FailSafe pipeline can automatically generate failure-recovery data at scale, adapt across multiple tasks and VLA models, and boost both the accuracy and robustness of robotic manipulation with the generalization capability across different spatial, viewpoints and embodiments.
IV-C Qualitative Analysis
@@ V Conclusions
In this paper, we propose FailSafe, a novel pipeline automatically generates failure reasoning data with executable recovery actions. FailSafe can be applied to any task in any simulator with motion planning functionality, enabling scalable failure dataset generation. Our experiments show that FailSafe-VLM exhibits strong failure reasoning capability and outperform other state-of-the-art VLMs. We further integrate FailSafe-VLM into VLAs to show its capability to recover from failures and boost overall performance in robotic manipulation tasks. FailSafe-VLM could also generalize cross spatial setups, camera views and robot embodiments. We hope that FailSafe pipeline will benefit the community by enhancing the essential failure reasoning capability of current robot control methods, paving the way toward more robust and explainable embodied AI applications.

### v2

@@ Abstract
Recent advances in robotic manipulation have integrated low-level robotic control into Vision-Language Models (VLMs), extending them into Vision-Language-Action (VLA) models. Although state-of-the-art VLAs achieve strong performance in downstream robotic applications, supported by large-scale crowd-sourced robot training data, they still inevitably encounter failures during execution. Enabling robots to reason and recover from unpredictable and abrupt failures remains a critical challenge. Existing robotic manipulation datasets, collected in either simulation or the real world, primarily provide only ground-truth trajectories, leaving robots unable to recover once failures occur. Moreover, the few datasets that address failure detection typically offer only textual explanations, which are difficult to utilize directly in VLA models. To address this gap, we introduce FailSafe, a novel failure generation and recovery system that automatically produces diverse failure cases paired with executable recovery actions. FailSafe can be seamlessly applied to any manipulation task in any simulator, enabling scalable creation of failure–action data. To demonstrate its effectiveness, we fine-tune LLaVa-OneVision-7B (LLaVa-OV-7B) to build FailSafe-VLM. Experimental results show that FailSafe-VLM successfully helps robotic arms detect and recover from potential failures, improving the performance of three state-of-the-art VLA models ($\pi_{o}$-FAST, OpenVLA, OpenVLA-OFT) by up to 22.6% on average across several tasks in Maniskill. Furthermore, FailSafe-VLM could generalize across different spatial configurations, camera viewpoints, object and robotic embodiments. We plan to release the FailSafe code to the community. Project Page: https://jimntu.github.io/FailSafe/
@@ I INTRODUCTION
Figure 1: An illustration of the FailSafe pipeline generating failure scenarios and corresponding executable recovery actions (above). Leveraging these, FailSafe enables FailSafe-VLM (below) to detect and recover from robot failures, while generalizing across different spatial configurations, viewing angles, object and embodiments.
@@
Furthermore, unlike prior work that frames failure reasoning merely as textual explanations [22, 19, 20], FailSafe collects executable recovery actions that can be directly applied by robots in real time. To ensure accuracy and robustness, each recovery action is validated through a rigorous systematic verification, confirming that it effectively resolves the failure scenario and eventually leads to successful task completion. By incorporating multi-view images and task instructions, this pipeline produces a customized, accurate, and explainable FailSafe dataset for robotic manipulation that would be prohibitively difficult to obtain through real-world data collection.
Experimental results further highlight the effectiveness of FailSafe in VLA deployments. We fine-tune LLaVa-OV-7B [23] to obtain FailSafe-VLM, which demonstrates strong failure reasoning capabilities that generalize across diverse spatial configurations, camera viewpoints, different objects and robotic embodiments. FailSafe-VLM significantly outperforms state-of-the-art VLMs, such as GPT-4o [24] and Gemini-2.5-flash [25], in detecting failures and generating accurate recovery actions on unseen failure trajectories. Furthermore, when serving as an external assistant to VLA models, FailSafe-VLM identifies and corrects potential failures in real time during robot execution, leading to an average performance improvement of up to 22.6% on ManiSkill tasks [26] compared to three VLA baselines without FailSafe-VLM. Finally, FailSafe-VLM also proves effective when evaluated on a previously unseen robot arm, demonstrating its ability to generalize across different embodiments.
@@
  • We show that FailSafe dataset enables existing VLMs to reason about failures and significantly improves the performance of VLA models across diverse viewpoints, spatial configurations, object categories and robotic embodiments.
@@ II-A Vision-Language-Action model
Figure 2: Top: Overall pipeline of FailSafe, which includes the autonomous generation of failure trajectories (I) and collection of delta recovery action (II). Failure-Action data pairs are passed to the next step only after a systematic verification (III) ensures the effectiveness of recovery action. Bottom: The FailSafe dataset (IV) is then used to fine-tune FailSafe-VLM, which is able to help robotic arms recover from failure cases (V).
@@ III FailSafe
We propose a novel pipeline, FailSafe, to systematically collect failure cases and corresponding recovery actions within ManiSkill [26]. While developed in ManiSkill, this pipeline is designed to generalize across a broader range of simulators that support motion planning. To ensure the reliability of the collected data, we incorporate a systematic verification mechanism that validates each failure–action pair before inclusion in the dataset. This step is crucial to guarantee both the diversity and effectiveness of the recovery actions. Details of the failure generation, action collection and systematic verification mechanism are provided in Section III-A, Section III-B and Section III-C, respectively.
FailSafe dataset is subsequently used to fine-tune LLaVa-OV-7B [23], yielding FailSafe-VLM, which is an expert vision–language model capable of assisting VLA models in reasoning and recovering from failures during robotic manipulation. The dataset format and the fine-tuning process for obtaining FailSafe-VLM are described in Section III-D and Section III-E, respectively.
@@ III-B Action Collection
III-C Systematic Verification
After getting the candidate corrective actions, they need to go through rigorous systematic verification to be included into the final FailSafe dataset. Importantly, the systematic verification is designed to ensure that the collected recovery actions $\Delta A$ are both robust in correcting failure cases and effective when applied at any timestep where failures could potentially occur in the immediate future.
In systematic verification, FailSafe replays the trajectory by incorporating the two poses ($P_{d}$ and $P_{c}$) collected earlier. Specifically, the motion planner first moves the robotic arm to the deviated pose $P_{d}$, then to the corrective pose $P_{c}$, and finally continues with the subsequent poses of the task (i.e., $A\rightarrow P_{d}\rightarrow P_{c}\rightarrow B\rightarrow C\rightarrow D$). If the originally failed operation is successfully completed after applying the corrective pose $P_{c}$, the corrective action $\Delta A$ along with the information such as image observation and failure types collected during the failure generation phase is added to the dataset. This process ensures both the reliability and diversity of the collected recovery actions, thereby enhancing FailSafe-VLM’s ability to help VLA models recover from failure scenarios in manipulation tasks.
@@ IV Experiments
In this section, we design a series of experiments to evaluate the failure reasoning and recovery capabilities of FailSafe-VLM in Framka Emika Panda robot arm. First, we roll out test seeds in Maniskill to generate unseen failure and success trajectories with new spatial configurations, and compare FailSafe-VLM with three other VLMs to assess the accuracy of predicted failure types and recovery actions. Second, we integrate FailSafe-VLM with three state-of-the-art VLA models to evaluate whether it can help them recover from failures and improve overall performance. In addition, we further investigate the generalization capability of FailSafe-VLM across different objects and robotic embodiments. Finally, a qualitative analysis is conducted to visualize the end-effector pose when the VLA model is integrated with FailSafe-VLM.
@@ IV-B Performance on VLA
TABLE III: Success rates of VLA models with and without FailSafe-VLM on three ManiSkill tasks using the Franka Emika Panda arm. Improvements are reported under the setting where the experimental camera view matches the VLA training view but is novel to FailSafe-VLM.
The experimental results in Table III show that all three VLA models yield average performance improvements with the inclusion of FailSafe-VLM. In particular, OpenVLA improves by more than 20%, with an average gain of 22.6% across the three tasks. For OpenVLA-OFT and $\pi_{o}$-FAST, whose baselines are already relatively strong, failure reasoning still pushes the boundary, increasing success rates by 8.0% and 4.0%, respectively.
For seeds recovered by FailSafe-VLM, we observe helpful nudges when the robotic arm becomes stuck or is about to fail. We would like to highlight that FailSafe-VLM should be viewed as an assistant for failure reasoning and recovery, rather than as a replacement for VLA models. Overall, these results demonstrate that the FailSafe pipeline can automatically generate failure-recovery data at scale, adapt across multiple tasks and VLA models, and boost both the accuracy and robustness of robotic manipulation.
IV-C Generalization Capability
To evaluate the generalization capability of FailSafe-VLM, we modify the objects used in three manipulation tasks. Specifically, Sphere and Charger are selected as novel object categories that FailSafe-VLM has never encountered during training. As shown in Table IV, FailSafe-VLM successfully assists the VLA model in correcting failures even with these unseen objects, achieving an average improvement of 17.4% across the three new tasks. This generalization ability can be attributed to the fundamental similarity of failure scenarios, as failure modes across different manipulation tasks often exhibit common patterns. The FailSafe pipeline enables the model to internalize these underlying failure principles, thereby empowering it to reason and recover from failures involving unseen objects in similar task settings.
We further evaluate whether FailSafe-VLM can assist VLA models in an embodiment unseen during training. Similarly, we also collect 1,000 trajectories per task on xArm 6 to fine-tune OpenVLA-OFT, while reusing the same FailSafe-VLM checkpoint trained exclusively on Franka Emika Panda robots. Although FailSafe-VLM is never trained on xArm 6 data, the failure scenarios and trajectories generated by FailSafe are independent of the specific robotic embodiment. Consequently, as shown in Table V, FailSafe-VLM exhibits cross-embodiment generalization by boosting the performance of the stack cube task on xArm 6 from 56% to 76% without degrading performance on other tasks.
VLA model FailSafe- VLM Pick Sphere Place Sphere Pick Charger Average
OpenVLA-OFT [3] 44.0% 36.0% 80.0% 53.3%
68.0% 52.0% 92.0% 70.7%
$\Delta$ +24.0% +16.0% +12.0% +17.4%
TABLE IV: Success rates of OpenVLA-OFT with and without FailSafe-VLM for novel object categories.
VLA model FailSafe- VLM Pick Cube Push Cube Stack Cube Average
@@ IV-C Generalization Capability
TABLE V: Success rates of OpenVLA-OFT with and without FailSafe-VLM for novel robotic embodiment.
IV-D Qualitative Analysis
@@ V Conclusions
In this paper, we propose FailSafe, a novel pipeline automatically generates failure reasoning data with executable recovery actions. FailSafe can be applied to any task in any simulator with motion planning functionality, enabling scalable failure dataset generation. Our experiments show that FailSafe-VLM exhibits strong failure reasoning capability and outperform other state-of-the-art VLMs. We further integrate FailSafe-VLM into VLAs to show its capability to recover from failures and boost overall performance in robotic manipulation tasks. FailSafe-VLM could also generalize cross spatial setups, camera views, object categories and robot embodiments. We hope that FailSafe pipeline will benefit the community by enhancing the essential failure reasoning capability of current robot control methods, paving the way toward more robust and explainable embodied AI applications.